Nick Brewer: Difference between revisions

Jump to navigation Jump to search
no edit summary
No edit summary
No edit summary
'''<big>7/22/14</big>'''
We tried to see some absorption again today. The data is [https://wiki.physics.wisc.edu/yavuz/images/4/47/Absorption_Data_2014-7-21.pdf here], and looks unconvincing. There is not a strong peak at 1055.076 nm like we are expecting. I think in order to get a better spectrum we will need to measure fluorescence from the crystal like in [http://www.opticsinfobase.org/DirectPDFAccess/B2A160D3-AED0-E02F-14B040E3D6801A62_7350/josab-11-4-591.pdf?da=1&id=7350&seq=0&mobile=no Shen's paper] instead of trying to directly measure absorption. [[Zjs log|Zach]] has been working on the locking circuit and we are now able to get a much less noisy signal. The fast feedback is fed directly to the diode instead of going through the current driver. Before the signal fluctuated ~25% or more when locked but now it is much less and we are able to get a clean signal out. There are still certain wavelengths where the error signal looks more like a peak than a dispersion graph and we are unable to lock, but we haven't yet determined the cause..
 
 

Navigation menu